OpenStudy (anonymous):

URGENT PLEASE: given that sin 2pi/7 = 0.7818 evaluate TO 4 decimal places: A) cos (3pi/14) B) cos (11pi/14)

OpenStudy (anonymous):

using equivalent algebraic expressions

OpenStudy (turingtest):

pulling out trig cheat sheet...

OpenStudy (anonymous):

ok lol...I will wait for you :D

OpenStudy (anonymous):

i just need the concept of how you go about doing these kinda questions...

OpenStudy (turingtest):

actually I have to go do what I would do: look at this sheet for some relevant formulae http://tutorial.math.lamar.edu/pdf/Trig_Cheat_Sheet.pdf

OpenStudy (turingtest):

make use of things like half-angle formulas, things like \(\sin(\frac\pi2-\theta)=\cos\theta\)

OpenStudy (turingtest):

looks like you almost have \(\large \frac{2\pi}7=\frac{4\pi}{14}=\frac{\pi}{14}+\frac{3\pi}{14}\) so maybe some of those addition/subtraction formulas will come in handy

OpenStudy (anonymous):

ok I get how I can manipulate the term in the middle of what is given for example in this care the 2pi/7 but how do I go from there to finding the cos 3pi/14 term

OpenStudy (anonymous):

like I can say: \[\sin(\pi/14 +3\pi/14) = \sin(\pi/14)\cos(3\pi/14) +\cos(\pi/14)\sin(3\pi/14) \] but I don't know if that will help

OpenStudy (turingtest):

you're basically trying the same type of stuff that I would, so continue looking at double-angle formulas to get that pi/14 argument to be a pi/7 argument I'm sorry I have to go, sorry to bail this time :( good luck!

OpenStudy (anonymous):

lol its iight thanks a lot man, I like the cheat sheet btw

OpenStudy (anonymous):

k wait i gotta read the question

OpenStudy (anonymous):

haha i dont know what to do lol

OpenStudy (anonymous):

its i got this one: so the solution is, for A) you say since \[\sin(\theta)=\cos({\pi/2} - \theta)\] then sin(4pi/14)= (cos pi/2 - 4pi/14) = cos (3pi/14) = 0.7818 for B) you use sin(theta) = -cos(pi/2+theta) and answer is= -0.7818

OpenStudy (anonymous):

ohhh i see

OpenStudy (anonymous):

haha i never done this stuff lol

OpenStudy (anonymous):

it looks fun

OpenStudy (anonymous):

yeh it is fun... I have done it 2 years ago, I cant remember anything now :S

OpenStudy (anonymous):

haha i have never done this but like i wanna learn it but i am dead at this point. So i cant really think

OpenStudy (anonymous):

i am actually gonna save that trig sheet on my comp. It is a good thing to have

OpenStudy (anonymous):

yeh it is very good i love it :D

OpenStudy (anonymous):

so who r u tutoring?

OpenStudy (anonymous):

some guy I don't know

OpenStudy (anonymous):

lol in ur neighbourhood?

OpenStudy (anonymous):

how do you tutor? give me tips and stuff

OpenStudy (anonymous):

kinda yeh

OpenStudy (anonymous):

well i am tutoring a grade 8er so its diff

OpenStudy (anonymous):

and do you ask them to tell you what they want you to teach them the day before?

OpenStudy (anonymous):

umm i dont but if its hs math then yes u do

OpenStudy (anonymous):

like u need to prepare

OpenStudy (anonymous):

but like u explain the topic and then u do one example together and then u let the guy do another one on his own

OpenStudy (anonymous):

and if he needs help then help him out on the way but try not to give away the answer

OpenStudy (anonymous):

ok cool i get that, now lets say you get across a question that you don't know lol

OpenStudy (anonymous):

what do you do in that case?

OpenStudy (anonymous):

ummmm i just say i dont know but u shld say like i dont know this let me work this out and get back to u tom. ok like even if u dont tutor him th enext day like email him the answer

OpenStudy (anonymous):

and if this guy is academically low then sometimes explaining the logic of a concept can confuse him so u just explain him the rules. It all depends on the student

OpenStudy (anonymous):

i see.. yeh yeh i get what you mean

OpenStudy (anonymous):

ok samjordan, amma go sleep am so tired

OpenStudy (anonymous):

gd night.

OpenStudy (anonymous):

ok bye